# A Bayesian framework for active object recognition, pose estimation and shape transfer learning through touch

### v1

@@ IV-B2 Sampling efficiency of the particle filter
To demonstrate the sampling efficiency of our particle filter, we tracked the number of particles during the GPIS-DHD object recognition and pose estimation experiments. The largest number of particles observed at any time step over 100 runs was 6917, for representing the distribution of 10 object classes and the pose of the object.
@@ IV-B3 Shape reconstruction with unknown objects
TABLE II: Comparison of two shape reconstruction methods
TWD/reconstruction error
MLE shape GPIS with MLE prior Poisson surface
class mean std mean std mean std
@@
Table II summarizes the statistics of the reconstruction error for each class. Each column reports the statistics of the TWD between the corresponding method and the ground truth shape. MLE shape denotes using the MLE prior from PF as the reconstructed shape, GPIS with MLE prior denotes our method, and Poisson surface denotes the Screened Poisson reconstruction method[31] (depth = $9$). As ice cream 2 and guitar 2 were regarded as known object by the framework, their GPIS results were not available, denoted with N/A. The same data were used for the GPIS and the Poisson surface reconstruction, which are the oriented point clouds acquired from the GPIS-DHD exploration experiments. In our experiment, the GPIS with MLE prior yields, in most cases, lower reconstruction error than the Poisson surface reconstruction.
Figure 9: Examples of the maximum likelihood estimation (MLE) priors and reconstructed surfaces in the experiments. Each row is one experiment. From left to right, the first column shows the ground truth object shape, the second column shows the MLE shape from the particle filter (PF), the third column shows the Gaussian process implicit surface (GPIS) reconstructed using the MLE shape as prior, the last column shows the reconstructed result using the Screened Poisson surface reconstruction.
A few examples of GPIS shape reconstruction with MLE shape priors can be seen in Fig 9. As can be seen, the proposed framework made sensible choices of the MLE priors and the GPIS reconstruction program did well in closing the gap between the MLE priors and the ground truth shapes.
@@ V-A Object recognition and pose estimation for known objects
In this work, we presented a Bayesian framework for robotic arm control that can efficiently identify objects from their shape and poses. Moreover, we showed knowledge transfer from known shapes to actively learn novel objects’ shapes of previously unfamiliar objects. Accordingly, adding the newly learned shape of the novel object lead to the prior being extended with a new object class that made the robotic arm identically successful in recognising objects in consequent trials. In order to do so, we formulated object recognition, shape and pose estimation with touch as a non-linear Bayesian filtering problem and adopted a particle filter-based approach (PF).
@@
Compared to Kalman Filters, PFs can approximate arbitrary distributions, which is needed to capture the ambiguity and uncertainty present due to partial tactile observations. Our PF uses a sampling procedure based on point-pair features and a custom weight assigning scheme to achieve efficient estimations of the joint distribution of object class and object pose estimation. We estimated the joint distribution of object class and object pose for ten object classes with less than 7000 particles, thus 700 particles on average per object. It is worth pointing out that the core idea behind the sampling procedure is to use translation and rotation invariant features, in our case, distances and angles, to find possible point correspondences, and then use the point correspondences to solve for the potential object poses, making it far more efficient than considering single points. This idea is not limited to the point-pair feature. For instance, three or four points could be considered simultaneously. Nonetheless, a set of two points is chosen in this study because the alignment of two oriented point pairs can uniquely determine the pose while the number of all point pair combinations is tractable. If more than two points per set are considered simultaneously, the number of possible combinations quickly becomes intractable. Similarly, in the sampling scheme, each object was represented by 200 points to keep the number of combinations tractable. Increasing the number of points will likely make the framework achieve the desired pose error sooner but at the cost of more computational resources.
@@
In the object recognition and pose estimation experiment with known objects, the framework recognized the correct object class and achieved desired pose estimation error with a small amount of data points regardless of the selected exploration procedure. Both the GPIS-DHD and RRT-based exploration procedure worked well, however GPIS-DHD generally covered the estimated surface faster than the RRT-based exploration procedure. For most known objects used in the experiment, data points covering only a fraction of the object surface were sufficient to infer the correct object class and object pose. The exception was the mug shown in Fig. 8, for which the rotational symmetry of its main body caused the pose to only be resolved by finding its handle. Compared to the RRT-based exploration procedure, the GPIS-DHD exploration procedure tries to actively touch the region on the estimated shape that is far away from existing data points. When the mug’s main body is covered by some data points, the handle of the estimated mug will be targeted more quickly. By confirming or failing to establish contact at the target location, the framework updates its belief of the handle location. In this case, the GPIS-DHD exploration procedure generated no-contact points during exploration, which was not the case for the RRT-based exploration procedure, since moving along the surface in small steps from a point on the surface guaranteed contact. The non-contact points in the GPIS-DHD played a crucial role in finding the mug’s handle, as it provided information to narrow down where the handle was located. The proposed DHD termination made sure that the exploration would continue until the handle was found, as failing to establish contact on the handle would result in a large DHD from the estimated surface to data points, thereby failing to fulfill the termination criterion.
@@ V-B Shape reconstruction for novel objects
The PF can identify unknown objects but lacks the ability to learn new shapes, so the GPIS is introduced to perform shape reconstruction. Aiming to exploit the prior knowledge, in this study, we initiate the prior of the GPIS to be the MLE shape from the PF. Our results showed that GPIS with MLE priors worked effectively with the novel object set and it achieved better performance than the screened Poisson surface reconstruction with the sparse point clouds from the GPIS-DHD exploration procedure. However, as the number of data points becomes denser, we expect the Poisson surface reconstruction may perform better than the GPIS with MLE priors, as the advantage of having a good prior decays as more data points are collected. In addition, dense point clouds can bring the covariance matrix (thin-plate kernel) of the GPIS closer to singularity and can potentially lead to broken surfaces. When observing Figs. 6d and 6h, it becomes clear that, in general, the framework took longer to explore a novel object than a known object. That is partly due to the greater applicability of the prior knowledge for known objects than for novel objects. In fact, if the object does not deviate enough from all known objects, it would be recognized as a known object. As was the case for Ice cream 2 and guitar 2 in our set. The active tactile exploration for these two objects were merely slightly slower than the known Ice cream 1 and Guitar 1 respectively. We assumed the success of the GPIS with sparse data stemmed from good priors, and we expected the GPIS to under-perform when no good priors were available. Surprisingly, in some cases, e.g., the dragon class in Table II, showed that despite the large discrepancy between the MLE prior and the ground truth, the GPIS still performed well and achieved low reconstruction error. One potential reasoning behind this observation is that even if the MLE shape is not a good fit globally, parts of it could have similar features to the actual shape. The locally similar parts were kept, and taken advantage of, whereas the mismatched parts were corrected by the GPIS.
@@ V-C Computational cost
When it comes to computational complexity, the PF’s computational cost scales linearly with the number of observations, because the PF’s weight assignment scheme for newly sampled particles requires the evaluation of one particle per class on all previous observations. Since the prior of the GPIS can be different at each time step, sparse GP with inducing points[36, 37] is not directly applicable, and the GPIS becomes computationally expensive as more data points are collected due to its $O(N^{3})$ complexity. One potential way to reduce the computational cost of the GPIS is to fix the prior after a certain amount of data, or when a certain level of surface coverage is reached, such that sparse GP becomes applicable. Moreover, as shown in the results section, new shapes can be learned and added as new priors to the framework, however, it will require proportionally more memory and computational time as the number of priors (here shape primitives) increases. Addressing these computational challenges would be one direction of future work. In this paper, the disadvantage of the increasing computational cost over time was partially compensated by the use of the proposed exploration procedure and the termination criterion. Noteworthy, though the proposed DHD threshold termination criterion was used in our experiments, it is by no means compulsory. A low DHD threshold can lead to a larger number of data points with better coverage of the object surface, while a high DHD threshold can allow faster conclusions but with higher chances of errors. A trade-off has to be made to determine a proper threshold value. However, the DHD from the estimated shape to existing data points could be viewed as a measure for the lowest data point density on the estimated object surface. In the case where another termination criterion applies, the DHD would still provide valuable information about how well the surface is covered, and the PF and GPIS can still output meaningful distribution on the object class, object pose, and estimated shape. Being uncertainty-aware of the coverage of the surface is one of the advantages of the proposed Bayesian framework.
@@ V-D Limitations and future work
As can be seen in the results of GPIS reconstruction result, there were defects such as small holes on thin surfaces and disconnected small parts. These defects were not repaired in the current framework, and this was reflected in the reconstruction error. Addressing these undesired defects from the learned models will be valuable future work. In addition, as the framework assumes no movement nor multiple objects in the scene, lifting these assumptions will be another potential work direction. Moreover, in this study, the framework determined if an object was known or novel based on the proposed MLE model evidence from the PF. As this threshold for the decision can vary, it provides some flexibility in choosing how similar the object should be to be classified as an known object. Ideally, classification should also take into account the topology of shapes, therefore this could also be a topic for future investigation. Last but not least, the experiments of this study were only performed in simulation as a proof of concept and to introduce the framework, however, validation with a robot hardware setup is in preparation and the results will be shown in a follow-up paper.
@@ VI Conclusion
In this paper, we proposed a Bayesian framework for active tactile object recognition, pose estimation, and shape reconstruction for motion planning of a robotic arm. The framework can utilize its prior knowledge to explore, recognize and localize known objects, and it can transfer the prior knowledge to learn shapes of novel objects. A GPIS-DHD exploration procedure and a DHD-based termination criterion were proposed to guide the active data acquisition process with success. In the experiments, we showed that the proposed framework along with the proposed exploration procedure were effective in the task of object recognition, pose estimation and shape reconstruction for both known and novel objects. To further demonstrate the learning capability of the framework, a newly learned object shape was added to the prior of the framework as a new object class and followingly, the framework successfully recognized the new object that it had learned from a previous trial through active tactile exploration.
Acknowledgment

### v3

@@ IV-B2 Sampling Efficiency of the Particle Filter
To demonstrate the sampling efficiency of our particle filter, we tracked the number of particles during the GPIS-DHD object recognition and pose estimation experiments. The largest number of particles observed at any time step over 100 runs was 6914, for representing the distribution of 10 object classes and the pose of the object.
@@ IV-B3 Shape Reconstruction with Unknown Objects
TABLE II: Comparison of shape reconstruction methods
Reconstruction error (TWD)
MAP shape GPIS with MAP prior Screened Poisson
class mean SD mean SD mean SD
@@
Table II summarizes the statistics of the reconstruction error for each class. Each column reports the statistics of the TWD between the corresponding method and the ground truth shape. MAP shape denotes using the MAP prior from the PF as the reconstructed shape, GPIS with MAP prior denotes our method, and Poisson surface denotes the Screened Poisson reconstruction method[29] (depth = $9$). As ice cream 2 and guitar 2 were regarded as known objects by the framework, their GPIS results were not available, denoted with N/A. The same data were used for the GPIS and the Poisson surface reconstruction, which are the oriented point clouds acquired from the GPIS-DHD exploration experiments. In our experiment, the GPIS with MAP prior yields, in most cases, lower reconstruction error than the Poisson surface reconstruction.
Figure 9: Examples of the maximum a posteriori (MAP) priors and reconstructed surfaces in the experiments. Each row is one experiment. From left to right, the first column shows the ground truth object shape, the second column shows the MAP shape from the particle filter (PF), the third column shows the Gaussian process implicit surface (GPIS) reconstructed using the MAP shape as prior, the last column shows the reconstructed result using the Screened Poisson surface reconstruction.
A few examples of GPIS shape reconstruction with MAP shape priors can be seen in Fig 9. As can be seen, the proposed framework made sensible choices of the MAP priors and the GPIS reconstruction program did well in closing the gap between the MAP priors and the ground truth shapes.
@@ V-A Object Recognition and Pose Estimation for Known Objects
In this work, we presented a Bayesian framework that can efficiently identify the object’s class and pose with the capability of transferring knowledge from known shapes to learn novel objects’ shapes through active touch. Accordingly, adding the newly learned shape of the novel object lead to the prior being extended with a new object class that made the framework identically successful in recognizing objects in consequent trials. In order to do so, we formulated object recognition, shape and pose estimation with touch as a non-linear Bayesian filtering problem and adopted a particle filter-based approach (PF).
@@
Compared to Kalman Filters, PFs can approximate arbitrary distributions, which is needed to capture the ambiguity and uncertainty present due to partial tactile observations. Our PF uses a sampling procedure based on point-pair features and a custom weight assigning scheme to achieve efficient estimations of the joint distribution of object class and object pose. At each time step, we approximated the joint distribution of object class and object pose for ten object classes with less than 7000 particles, thus 700 particles on average per object. It is worth pointing out that the core idea behind the sampling procedure is to use translation and rotation invariant features, in our case, distances and angles, to find possible point correspondences, and then use the point correspondences to solve for the potential object poses, making it far more efficient than considering single points. This idea is not limited to the point-pair feature. For instance, three or four points could be considered simultaneously. Nonetheless, a set of two points is chosen in this study because the alignment of two oriented point pairs can uniquely determine the pose while the number of all point pair combinations is tractable. If more than two points per set are considered simultaneously, the number of possible combinations quickly becomes intractable. Similarly, in the sampling scheme, each object was represented by 200 points to keep the number of combinations tractable. Increasing the number of points will likely make the framework achieve the desired pose error sooner but at the cost of more computational resources.
@@
In the object recognition and pose estimation experiment with known objects, the framework recognized the correct object class and achieved desired pose estimation error with a small amount of data points regardless of the selected exploration procedure. Both the GPIS-DHD and RRT-based exploration procedure worked well, however GPIS-DHD generally covered the estimated surface faster than the RRT-based exploration procedure. For most known objects used in the experiment, data points covering only a fraction of the object surface were sufficient to infer the correct object class and object pose. The exception was the mug shown in Fig. 8, for which the rotational symmetry of its main body caused the pose to only be resolved by finding its handle. Compared to the RRT-based exploration procedure, the GPIS-DHD exploration procedure tries to actively touch the region on the estimated shape that is far away from existing data points. When the mug’s main body is already covered by some data points, the handle of the estimated mug will be prioritized by the GPIS-DHD exploration procedure. By confirming or failing to establish contact at the target location, the framework updates its belief of the handle location. In this case, the GPIS-DHD exploration procedure generated no-contact points during exploration, which was not the case for the RRT-based exploration procedure, since moving along the surface in small steps from a point on the surface guaranteed contact. The non-contact points in the GPIS-DHD played a crucial role in finding the mug’s handle, as it provided information to narrow down where the handle was located. The proposed DHD termination made sure that the exploration would continue until the handle was found, as failing to establish contact on the handle would result in a large DHD from the estimated surface to data points, thereby failing to fulfill the termination criterion.
@@ V-B Shape Reconstruction for Novel Objects
The PF can identify novel objects but lacks the ability to learn new shapes, so the GPIS is introduced to perform shape reconstruction. Aiming to exploit the prior knowledge, in this study, we initiate the prior of the GPIS to be the MAP shape from the PF. Our results showed that GPIS with MAP priors worked effectively with the novel object set and it achieved better performance than the screened Poisson surface reconstruction with the sparse point clouds from the GPIS-DHD exploration procedure. However, as the number of data points becomes denser, we expect the Poisson surface reconstruction may perform better than the GPIS with MAP priors, as the advantage of having a good prior decays as more data points are collected. In addition, dense point clouds can bring the covariance matrix (thin-plate kernel) of the GPIS closer to singularity and can potentially lead to broken surfaces. When observing Figs. 6d and 6h, it becomes clear that, in general, the framework took longer to explore a novel object than a known object. That is partly due to the greater applicability of the prior knowledge for known objects than for novel objects. In fact, if the object does not deviate enough from all known objects, it would be recognized as a known object. As was the case for Ice cream 2 and guitar 2 in our set. The active tactile exploration for these two objects were merely slightly slower than the known Ice cream 1 and Guitar 1 respectively. We assumed the success of the GPIS with sparse data stemmed from good priors, and we expected the GPIS to under-perform when no good priors were available. Surprisingly, in some cases, e.g., the dragon class in Table II, showed that despite the large discrepancy between the MAP prior and the ground truth, the GPIS still performed well and achieved low reconstruction error. One potential reasoning behind this observation is that even if the MAP shape is not a good fit globally, parts of it could have similar features to the actual shape. The locally similar parts were kept, and taken advantage of, whereas the mismatched parts were corrected by the GPIS.
@@ V-C Computational Cost
When it comes to computational complexity, the PF’s computational cost scales linearly with the number of observations, because the PF’s weight assignment scheme for newly sampled particles requires the evaluation of one particle per class on all previous observations. Since the prior of the GPIS can be different at each time step, sparse GP with inducing points[34, 35] is not directly applicable, and the GPIS becomes computationally expensive as more data points are collected due to its $O(N^{3})$ complexity. One potential way to reduce the computational cost of the GPIS is to fix the prior after a certain amount of data, or when a certain level of surface coverage is reached, such that sparse GP becomes applicable. Moreover, as shown in the results section, new shapes can be learned and added as new priors to the framework, however, it will require proportionally more memory and computation time as the number of priors (here shape primitives) increases. Addressing these computational challenges would be one direction of future work. In this paper, the disadvantage of the increasing computational cost over time was partially compensated by using the proposed exploration procedure and the termination criterion. Noteworthy, though the proposed DHD threshold termination criterion was used in our experiments, it is by no means compulsory. A low DHD threshold can lead to more data points with better coverage of the object surface, while a high DHD threshold can allow faster conclusions but with higher chances of errors. A trade-off has to be made to determine a proper threshold value. The DHD from the estimated shape to existing contact points could also be viewed as a measure for the lowest data point density on the estimated object surface. In the case where another termination criterion applies, the DHD would still provide valuable information about how well the surface is covered, and the PF and GPIS can still output meaningful distribution on the object class, object pose, and estimated shape. Being uncertainty-aware of the coverage of the surface is one of the advantages of the proposed Bayesian framework.
@@ V-D Limitations and Future Work
As can be seen in the results of GPIS reconstruction result, there were defects such as small holes on thin surfaces and disconnected small parts. These defects were not repaired in the current framework, and this was reflected in the reconstruction error. Addressing these undesired defects from the learned models will be valuable future work. In addition, as the framework assumes no movement nor multiple objects in the scene, lifting these assumptions will be another potential work direction. Moreover, in this study, the framework determined if an object was known or novel based on the proposed MAP model evidence from the PF. As this threshold for the decision can vary, it provides some flexibility in choosing how similar the object should be to be classified as an known object. Ideally, classification should also take into account the topology of shapes, therefore this could also be a topic for future investigation. Last but not least, the experiments of this study were only performed in simulation as a proof of concept and to introduce the framework, however, validation with a robot hardware setup will be left as future work.
@@ VI Conclusion
In this paper, a unified Bayesian framework is proposed to address active tactile object recognition, pose estimation, and shape transfer learning. Through experiments in simulation, we demonstrated the proposed framework’s efficiency and effectiveness in tacking both known and novel objects. It can estimate the object class and object pose for known objects reliably and transfer the knowledge of known shapes to learn the shapes of novel objects. Furthermore, the learned shapes can be added to the framework’s prior knowledge for object recognition and pose estimation, enabling continual learning.
Acknowledgement
We would like to express our gratitude to Dr. Lorenzo Natale for his valuable feedback on the paper.